### 12 Apr 2026 v1

# Mapping High-Performance Regions in Battery Scheduling across Data Uncertainty, Battery Design, and Planning Horizons

Abstract
This study presents a triadic analysis of energy storage operation under multi-stage model predictive control, investigating the interplay between data characteristics, forecast uncertainty, planning horizon, and battery c-rate. Synthetic datasets are generated to systematically explore variations in data profiles and uncertainty, enabling parametrization and the construction of relationships that map these characteristics to optimal horizon length. Results reveal the presence of an effective horizon, defined as the look-ahead length beyond which additional forecast information provides limited operational benefit. Accounting for this horizon can reduce computational costs while maintaining optimal performance. The study provides optimal horizon lengths across a broad range of combinations of battery types, uncertainty levels, and data profiles, offering practical guidance for industrial storage operation. It also quantifies revenue losses due to forecast uncertainty, showing that errors can impact performance even for fast batteries. Finally, the framework lays the groundwork for future machine learning approaches that map dataset parametrization to optimal horizons, supporting continuous optimization in industrial settings without heavy computation.
1 Introduction
Battery energy storage systems play an increasingly central role in modern energy systems. They enable the integration of renewable generation, provide grid-balancing services, and constitute a profitable asset class through energy arbitrage and participation in electricity markets. Across these applications, a critical operational task is the scheduling of battery charge and discharge actions over time, typically based on forecasts of future electricity prices, load demand, or renewable generation.
Multi-stage battery scheduling problems are commonly formulated as optimization problems in which an objective function, such as operational cost or economic profit, is optimized subject to physical and operational constraints. For example, these constraints may include energy capacity, charge and discharge power limits, state-of-charge bounds, and so on Oudalov et al. (2007); Marwali et al. (1998). In its simplest form, this problem is often posed as a linear program, assuming constant efficiencies and power limits independent of the state of charge Park et al. (2017). More detailed formulations relax these assumptions and employ nonlinear optimization techniques, such as simulated annealing or other heuristic methods Ekren and Ekren (2010); Zhang et al. (2018); Maleki and Pourfayaz (2015).
Regardless of the specific formulation, battery scheduling inherently relies on forecasts of future values. As a result, uncertainty enters the optimization problem through forecast errors, which typically increase with the look-ahead horizon. Optimization under uncertainty is a long-established field, with roots dating back at least to the mid-20th century Dantzig (1955), and has gained renewed importance with the growing penetration of weather-dependent renewable energy sources Li and Grossmann (2021).
A wide range of methodological frameworks have been proposed to address forecast uncertainty, including stochastic optimization and dynamic programming Dantzig (1955); Bellman (1957); Pereira and Pinto (1991), robust optimization Soyster (1973), distributionally robust optimization Xu et al. (2012); Shafieezadeh-Abadeh et al. (2019), and reinforcement learning approaches Watkins (1989); Sutton (1988); Sui and Song (2020). Many of these sequential decision-making methods share a structural element similar to that used in control theory: a receding-horizon or rolling optimization framework, as formalized in Model Predictive Control (MPC) Richalet et al. (1978); Mayne et al. (2000). In MPC approaches, the optimization problem is repeatedly solved on a rolling basis as new information becomes available.
A fundamental design choice in any MPC framework is the length of the optimization window. If the horizon is too short, the optimization may become myopic and fail to anticipate upcoming opportunities or constraints. Conversely, if the horizon is too long, increasingly uncertain forecasts may degrade performance, such that incorporating additional future information becomes counterproductive. Additionally, the computation time required to obtain the solution to the optimization, is also driven by the length of this planning horizon. Therefore, for time-consuming optimization tasks executed in multi-stage setups, the use of the optimal horizon allows making informed decisions faster. This may be crucial when optimization outcomes are used in energy trading, where delayed decisions may result in significant financial losses. The choice of horizon length is therefore, a critical parameter in battery scheduling under uncertainty.
Despite its importance, the optimization window or horizon is often treated as a fixed or secondary design parameter. In many studies, the horizon length is calibrated for a specific application or market context Ofoe et al. (2025). Some works propose adaptive or dynamically varying horizons to better accommodate changing conditions Pous et al. (2025). In contrast, a large body of industry-driven studies report limited sensitivity to the chosen horizon length, and consequently adopt intuitive values (such as 24 hours) without detailed justification Weitzel and Glock (2018); Mercier et al. (2023); Diller et al. (2024).
This apparent insensitivity can be explained, at least in part, by a structural property of many battery systems used in practice. Batteries with high power-to-capacity ratios (high C-rates) can complete full charge–discharge cycles over relatively short time scales. As a result, even when a long optimization horizon is considered, the scheduling decisions that materially affect battery operation are predominantly influenced by near-term information. Far-future data, although formally included in the optimization problem, often has negligible impact on optimal decisions Zhong et al. (2023).
This observation motivates the concept of an effective planning horizon: the portion of the nominal optimization horizon over which future information can meaningfully influence scheduling decisions, given the battery’s physical constraints. Importantly, this effective planning horizon emerges endogenously from the interaction between battery dynamics and the temporal structure of the input data, and persists even under perfect foresight. In this sense, fast-cycling batteries exhibit a self-limiting effect on the optimization window, naturally reducing exposure to long-term forecast uncertainty .
In many practical settings, this self-limiting effect is further reinforced by the structure of forecast data. For example, electricity prices, photovoltaic generation, and wind power production are often driven by weather processes, for which short-term forecasts are relatively reliable. Consequently, fast lithium-ion batteries frequently make decisions before forecast uncertainty unfolds severely, which partially explains why horizon length may appear to have limited importance in industrial applications.
However, these conditions do not characterize all energy storage systems or market environments. There are several important cases in which the choice of optimization horizon remains critical. First, certain electricity markets, such as balancing or reserve markets, exhibit rapid growth of forecast uncertainty. This may render predictions unreliable over time spans as short as a few hours. Second, advanced methods for optimization under uncertainty, including stochastic optimization, often face steep computational complexity that increases rapidly with horizon length Fortenbacher et al. (2018). If satisfactory performance can be achieved with shorter horizons, this may enable the practical deployment of such methods, which are frequently dismissed in industrial contexts due to computational constraints Zhang et al. (2023); Xu et al. (2025).
Finally, lithium-ion batteries do not represent the full spectrum of energy storage technologies. Concerns regarding resource availability and sustainability have motivated increasing interest in alternative solutions, such as flow batteries and hydrogen-based energy storage systems, which are actively promoted in policy frameworks including those of the European Union Centre and Commission (2023). These technologies typically operate at significantly lower C-rates than lithium-ion batteries, resulting in slower charge and discharge cycles on the order of several hours. As a consequence, their effective planning horizons are substantially longer, and they are more exposed to the detrimental effects of forecast uncertainty over extended horizons.
In such systems, selecting an appropriate optimization horizon becomes essential to balance the value of longer-term information against the risks introduced by uncertainty. Understanding how this balance depends jointly on battery design and data characteristics is therefore critical for informed system design and operation.
Figure 1: Diagram illustrating the interplay between battery design parameters (e.g., C-rate and charge efficiency), data and uncertainty characteristics (underlying signal amplitude, frequency and phase, combined with stochastic forecast errors), and the planning horizon length used in the MPC optimization. Different combinations within this triangular space give rise to different profit levels, conceptualized here as an interior profit landscape.
This paper, provides a methodology to analyse the interplay between battery power–capacity ratios, data characteristics, and optimization horizon length (see Figure 1). The study is based on a deterministic multistage scheduling framework, to deliberately isolate the structural effects of horizon length without confounding influences from stochastic or risk-aware solution methods. Specifically, the methodology aims to (i) investigate how battery design influences the ability to exploit temporal variations in forecast data, (ii) examine how data properties (such as amplitude variability, seasonality patterns, and forecast error growth) affect the value of longer horizons, and (iii) identify high-performance regions in the combined space of battery characteristics, data profiles, and planning horizons. The resulting insights should provide guidance for selecting battery designs and optimization horizons adapted to the nature of the underlying data, and establish a foundation for future extensions incorporating stochastic optimization or other optimization under uncertainty approaches.
2 Literature review
The literature on uncertainty and MPC is extraordinarily vast. The origins of Model Predictive Control can be traced back to the work of Richalet et al. in the late 1970s Richalet et al. (1978), who introduced the term Model Predictive Heuristic Control (MPHC) in the context of digital control of multivariable industrial processes. The method, in its original form, was based on a continuous feedback loop in which a future system trajectory is being constantly predicted. Only the first control action is implemented before the system state is re-measured and the optimization repeated. This rolling procedure explicitly accounts for discrepancies between predicted and realized system behavior. The term heuristic reflected the use of predicted information and simplified optimization methods. While early MPC (MPHC) relied on heuristic model-based predictions with limited optimization guarantees, modern MPC Mayne et al. (2000) is formulated as an explicit constrained optimal control problem with well-defined objective functions and theoretical properties.
Meanwhile, the field of optimization under uncertainty saw very important developments much earlier than the first formulation of MPHC. Dantzig formulated linear programming under uncertainty in 1955 Dantzig (1955), introducing one of the first formal frameworks to optimize expected costs. Shortly after, Bellman extended his work on dynamic programming Bellman (1954) to multistage decision problems under stochastic uncertainty Bellman (1957), establishing a systematic approach to sequential optimization in uncertain environments. These foundational works laid the groundwork for modern stochastic and robust optimization Soyster (1973) methods. The scientific background driving the development of these methods included Control Theory and Automation and Control.
Despite these early formulations, it would take years if not decades, to see a substantial uptake of published applications in the energy sector (around the 90’s). This is especially true when considering applications to the scheduling of energy storage systems (e.g., a popular paper from 1995 Maly and Kwan (1995)). Later, with the explosion of renewable energies, research in these areas has seen a massive increase, including the development of new methods, such as distributionally robust optimization (DRO) Delage and Ye (2010); Xu et al. (2012) (highly inspired by a 1958 paper Scarf (1958)) and hybrid strategies combining these approaches Powell (2019); Bakker et al. (2020); Roald et al. (2023)). As a measure of relevance of these methods, a few years after its ‘modern’ formulation, DRO would go on to inspire ‘mass transportation’ regularization techniques in Machine Learning methods Shafieezadeh-Abadeh et al. (2019).
As mentioned in the introduction, the size of the MPC optimization window is a central theme in the present study. It is important to distinguish between minimal or sufficient horizons, which arise in deterministic settings where longer horizons do not degrade performance, and optimal horizons, which emerge in the presence of forecast uncertainty, where excessive look-ahead may be detrimental. In the present study, the focus is centered on the balance between myopic and farsighted decisions, or in other words, the optimal length. Several works have focused on determining the optimal window size under uncertain forecasts. And others, carry out a similar analysis in deterministic contexts without uncertainty, aiming to identify the minimal horizon. While the following references pertain to the pursuit of both minimal and optimal horizons, they are retained here to provide a comprehensive view of horizon-related studies. Within this classification, some of these studies propose dynamic or variable windows that adapt over time to better respond to important features in the data, such as daily seasonal patterns or extreme local behaviors Prat et al. (2024); Ofoe et al. (2025); Laguna et al. (2022); Bøhn et al. (2021); Hönen et al. (2023), while others assume a constant optimal window throughout the control horizon Kaiser et al. (2025); Mayhorn et al. (2017); Kannan and Zavala (2011); Houwing et al. (2007).
The latter studies (constant horizon) often focus on practical applications where the horizon is set based on engineering considerations or computational constraints. Works on storage and MPC sizing frequently adopt a fixed horizon defined by problem context or sampling choices, such as Kaiser et al. (2025), which analyzes how fixed horizon length affects energy absorption in wave energy converters’ MPC schedules. Mayhorn et al. and Kannan et al. Mayhorn et al. (2017); Kannan and Zavala (2011) investigate battery and microgrid storage scheduling under fixed horizons, demonstrating how horizon choice influences cost and reliability in deterministic optimization frameworks. Then, in Houwing et al. (2007), the work also examines fixed horizon MPC for control systems with conventional settings to balance performance and computational tractability. Collectively, these studies highlight that while the horizon may remain constant, its selection critically influences both solution quality and solver burden across diverse applications.
Among the former works (often referred to as adaptive horizon model predictive control or AHMPC), most papers in the literature are somewhat tied to particular domains or use-cases. But some others set out to develop more general approaches. In the first group, Bøhn et al. Bøhn et al. (2021) propose a reinforcement learning framework to learn the optimal horizon length as a function of state information, improving performance compared to fixed horizons. Then, in Laguna et al. (2022), the authors introduce a dynamic horizon selection methodology for building energy systems, demonstrating that horizon length can be adapted based on building thermal inertia and forecast models to reduce computation without sacrificing performance. Lastly, in Hönen et al. (2023), the authors propose a rolling-horizon optimization framework in which the prediction horizon is fixed, but the starting times of successive optimization problems are chosen dynamically. By adaptively deciding when to re-optimize, the method implicitly controls the effective use of future information and mitigates forecast uncertainty without modifying the nominal horizon length. These studies highlight that dynamically adapting the prediction horizon (both in length and resolution) can significantly enhance MPC performance while balancing computational effort.
In the second group, which is particularly relevant, Prat et al. Prat et al. (2024) introduce a framework to study finite-horizon energy storage scheduling problems, analyzing how forecast uncertainty and data characteristics influence horizon length. Ofoe et al. Ofoe et al. (2025) build on this discussion, proposing a method to detect when a finite horizon is sufficient in practice; their work explicitly references and discusses the contributions of Prat et al. Prat et al. (2024). Finally, in Alla et al. (2021), in contrast, present a time-adaptive MPC framework that adjusts both the prediction horizon length and the time step based on system dynamics, offering a general-purpose strategy to balance computational efficiency and control accuracy. The strategy for this dynamic adjustment consists of using an error estimate to apply finer steps where the system changes rapidly and coarser steps where it evolves slowly, improving efficiency while maintaining accuracy.
Another relevant aspect worth noting is that, although motivated by uncertainty, the present study proposes a ‘rigid’ horizon under a linear and deterministic framework. This choice provides a neutral ground to systematically compare the interplay between different data profiles and battery characteristics. It should be clarified, however, that otherwise, approaches often account for uncertainty explicitly by treating the end of the optimization window not as a hard limit, but as a soft or fading horizon. Strategies of this kind, such as horizon discounting, are widespread in the literature across various optimization frameworks, including MPC Schwenkel et al. (2024, 2025), dynamic programming for Markov decision processes, Ilhuicatzi-Roldán et al. (2017), Stochastic Optimization (common since the earliest formulations Bellman (1957); Scarf (1958)), Robust Optimization Nilim and El Ghaoui (2005) or chance constraints Yan et al. (2018), among other approaches, such as hybrid robust-stochastic methods Brüdigam et al. (2020).
Finally, methods for handling uncertainty within a fixed predictive horizon are abundant. For example, there are robust or probabilistic MPC for microgrids that explicitly balance forecast error decay against scheduling performance Aguilar et al. (2024) or comprehensive surveys of MPC‑based energy management strategies Ordóñez et al. (2024). However, few studies systematically address how to choose or adapt the horizon length itself in the presence of uncertainty across battery systems and input data characteristics. Works like Prat et al. Prat et al. (2024) explore principled finite‑horizon approximations for energy storage scheduling, analyzing how data profiles and forecast error growth constrain the effective look‑ahead, while related real‑time scheduling studies discuss minimum sufficient horizons in the smart grid context Ofoe et al. (2025). By contrast, the literature on shrinking‑horizon predictive control, such as frameworks that progressively shorten the horizon while solving the scheduling optimization to balance cost and forecast reliability, illustrates alternative strategies to modulate the planning window itself Caminiti et al. (2024). Despite these innovative techniques, a general exploration that connects horizon length with uncertainty characteristics, forecast dynamics, and storage design remains underdeveloped. Such exploration is the main contribution of the present paper.
3 Methods
This study proposes a methodology to systematically explore how different data characteristics and battery specifications influence the optimal planning horizon in multi-stage scheduling optimization. Because real-world data profiles vary widely, they are difficult to compare directly. For instance, these may range from electricity prices in different markets to renewable generation patterns and consumption from retail or industrial loads. To address this issue, data profiles are constructed through a parameterized model. By representing each dataset through a set of meaningful parameters, datasets can be compared in a common parametric space and be systematically studied according to their influence on optimization outcomes.
Once the parametric model is established, it will be possible to evaluate its representational fidelity using standard metrics such as mean absolute error (MAE) and mean squared error (MSE), providing readers with an intuitive sense of how closely the model approximates real data. This model will serve as a generative tool to produce a wide spectrum of datasets, encompassing both ground truth and forecast data. The profiles begin with simple, theoretical proof-of-concept datasets to highlight fundamental behaviors of the system and then move toward more complex datasets that closely resemble real-world scenarios.
With these datasets in hand, a multi-stage model predictive control (rolling horizon linear optimization) is defined using a grid search approach to identify optimal planning horizons. This procedure will allow for the systematical mapping of the relationship between dataset characteristics (including forecast uncertainty) and optimal horizon windows for a preselected set of battery types. These will range from fast-responding lithium batteries, to slower technologies such as flow and hydrogen batteries. Alongside horizon mapping, the profitability associated with each battery type under different data conditions is also examined. This provides a comprehensive view of their performance under uncertainty.
Building on these results, in future work, it could be interesting to formalize a predictive function, $f(\cdot)$, that relates the properties of the battery and the parameters of the data (including uncertainty parameters in forecasts) to the optimal planning horizon. Special attention should be given to how well this function may be able to capture the influence of forecast uncertainty on the optimization window. The validity and practical relevance of $f(\cdot)$ would be assessed by comparing its predictions against real-world dataset scenarios, evaluating its potential as a tool for guiding preliminary design and operational strategies for battery-based energy storage systems. If successful, this approach would lead to cost-efficient dynamic horizon implementations, where the optimal length is be computed at every optimization step without computational burden.
3.1 Generation of synthetic data
3.1.1 Synthetic ground truth data
Ground-truth time series are generated using a hybrid parametric–stochastic construction that combines a Fourier-inspired deterministic component with a stochastic time-series model. The deterministic backbone is defined as a superposition of a finite number of sinusoidal components:
$x(t)=\sum_{k=1}^{K}A_{k}\sin\!\left(2\pi f_{k}t+\phi_{k}\right),$ (1)
where each component is characterized by an amplitude $A_{k}$, frequency $f_{k}$, and phase shift $\phi_{k}$. The time index $t$ is discretized uniformly with a fixed sampling step, yielding a sequence of length $n$.
This Fourier-inspired formulation is deliberately chosen for two reasons. First, it enables a bottom-up construction of synthetic time series with controllable temporal features such as multi-scale seasonality, periodic spikes, and smooth trends, which are commonly observed in energy-related data (e.g., prices, demand, and renewable generation). Second, it supports a top-down interpretation, as real-world datasets can be approximated or decomposed into dominant frequency components, allowing synthetic and empirical data to be represented in a common parametric space. This duality is essential for later mapping insights obtained from synthetic experiments to real datasets.
The raw Fourier signal in (1) is subsequently normalized and reshaped to match realistic magnitude constraints. Let $\tilde{x}(t)$ denote the normalized signal. A nonlinear shaping operation is applied:
$\hat{x}(t)=\mathrm{sign}(\tilde{x}(t))\,|\tilde{x}(t)|^{\gamma},$ (2)
where $\gamma>0$ controls the sharpness of peaks and troughs. Values $\gamma>1$ emphasize extreme events, while $\gamma<1$ produce flatter profiles. The shaped signal is then scaled to a predefined magnitude range.
To capture temporal dependencies and stochastic structure beyond purely periodic behavior, the deterministic component is combined with a stochastic process modeled using a seasonal autoregressive integrated moving average (SARIMA) model Box et al. (2015). Specifically, the final series is constructed as:
$y(t)=\hat{x}(t)+z(t),$ (3)
where $z(t)$ follows a SARIMA process that captures autocorrelation, persistence, and seasonal stochasticity. This component introduces realistic temporal dynamics such as clustered variability, lagged effects, and structured deviations from the deterministic backbone, which are not adequately represented by independent noise.
Compared to simple additive white noise, the SARIMA component allows for temporally correlated fluctuations and richer stochastic behavior, better reflecting real-world energy time series where shocks and deviations exhibit memory and seasonal structure. The parameters of the SARIMA model can be tuned to control the strength and persistence of these stochastic effects.
The final output may be clipped to enforce physical bounds (e.g., non-negativity for demand or price signals), and all random elements are generated using controlled seeds to ensure reproducibility.
Alternative approaches for synthetic data generation include state-space models Durbin and Koopman (2001); Kalman (1960), or data-driven generative models including variational autoencoders Kingma and Welling (2014); Fortuin et al. (2020) and generative adversarial networks Goodfellow et al. (2014); Yoon et al. (2019). While these methods can produce realistic samples, they typically offer less transparent parameterizations and reduced interpretability. In contrast, the hybrid approach adopted here preserves the intuitive, low-dimensional structure of Fourier-based representations while augmenting them with stochastic dynamics, enabling a balance between interpretability and realism. This is particularly important for establishing mappings between data characteristics, battery parameters, and optimal planning horizons.
3.1.2 Synthetic forecast data generation
Forecast time series are generated by perturbing the corresponding ground-truth series with stochastic errors, in order to emulate realistic prediction uncertainty. Concretely, for each time step $t$, the forecast value $y_{f}(t)$ is obtained as
$y_{f}(t)=y(t)+\epsilon(t),$ (4)
where $y(t)$ is the ground-truth series and $\epsilon(t)$ is a random error sampled from a normal distribution,
$\epsilon(t)\sim\mathcal{N}\bigl(\mu_{\epsilon}(t),\sigma_{\epsilon}(t)^{2}\bigr).$ (5)
The mean $\mu_{\epsilon}(t)$ is typically set to zero, representing unbiased forecasts, while the standard deviation $\sigma_{\epsilon}(t)$ captures the expected forecast uncertainty. In our setup, $\sigma_{\epsilon}(t)$ may evolve with the forecast horizon to reflect the intuitive notion that longer-term forecasts are generally less accurate. This evolution can follow different profiles—linear, exponential, or seasonally modulated—allowing flexible control over the growth of forecast errors over time.
To avoid unrealistic behavior arising from independent forecast errors, temporal dependence is enforced through an autocorrelated error process. Specifically, forecast errors are generated using a first-order autoregressive (AR(1)) model:
$\varepsilon_{t}=\rho\,\varepsilon_{t-1}+\sqrt{1-\rho^{2}}\,\eta_{t},\qquad\eta_{t}\sim\mathcal{N}(0,1),$ (6)
where $\rho\in(-1,1)$ controls the degree of autocorrelation. In practice though, only positive autocorrelation is considered in the present paper, reflecting the observed tendency of forecast errors to persist rather than oscillate. In any case, this formulation ensures that forecast errors exhibit persistence over time, reflecting the realistic behavior that prediction errors tend to cluster rather than occur independently. The resulting error process is subsequently scaled to match a prescribed variance profile and combined with the deterministic signal. This produces forecast trajectories in which deviations from the ground truth evolve smoothly and coherently over time, rather than as isolated random shocks. Compared to simple mean-reverting adjustments, this approach better captures the temporal structure of forecast uncertainty while remaining computationally lightweight and easily tunable. In future work, more sophisticated approaches may be considered, such as modeling full covariance structures for prediction errors Pinson et al. (2009).
Finally, the standard deviation of the error distributions is increased along the time axis to simulate the effect of degrading forecast (the longer the horizon, the less accurate the forecast). For this study, three methods have been implemented: linear, exponential and seasonal increase of the standard deviation.
Overall, this method provides a simple and interpretable framework for generating forecast series, complementing the Fourier-based ground-truth generation. The resulting synthetic forecasts provide a controlled way to study how forecast uncertainty propagates through the optimization process and affects the determination of optimal planning horizons. By varying the error magnitude and its evolution over the forecast horizon, one may systematically explore the sensitivity of different battery technologies to both the level and temporal structure of prediction errors. This makes the synthetic forecasts a valuable tool for linking data characteristics to operational performance in a manner that is both reproducible and interpretable.
3.2 Model Predictive Control schema
The methodology implements a rolling-horizon Model Predictive Control (MPC) scheme based on repeated linear programming optimizations, which are detailed in the next section. At each control step, the optimization algorithm computes an optimal sequence of control actions over a finite planning horizon, referred to here as the optimization window. Only a subset of these actions is executed before the optimization is repeated using updated information, following the standard MPC paradigm.
Figure 2 illustrates this process schematically for two consecutive optimization runs, denoted as optimization window 1 and optimization window 2. Each optimization window spans a fixed horizon length and takes as input the forecast data available at the time the optimization is launched. The forecast values used by the first and second optimization runs are denoted by $F_{1}$ and $F_{2}$, respectively, while the realized ground-truth values are denoted by $GT$.
Figure 2: Rolling-horizon Model Predictive Control (MPC) scheme with overlapping optimization windows. Two consecutive optimization runs (optimization window 1 and 2) are shown, using forecast inputs $F_{1}$ and $F_{2}$, respectively, and ground-truth values GTGT. The windows are shifted forward by a stride shorter than the optimization horizon, resulting in overlapping action sequences $\{x^{1}_{n}\}$ and $\{x^{2}_{n}\}$. Only the actions within the stride interval are retained (highlighted), while the remaining actions are discarded (light gray). The forecast publication interval is denoted by $(t_{F_{2}}-t_{F_{1}})$.
The time interval between successive forecast publications is denoted by $(t_{F_{2}}-t_{F_{1}})$. This interval determines when new forecast information becomes available and, consequently, when the optimization problem is re-solved. The distance between consecutive optimization windows is defined by the stride, which specifies how far the rolling window advances in time between optimization runs. Importantly, the stride may be smaller than the optimization window length, resulting in overlapping optimization horizons.
Because of this overlap, each optimization run produces a full sequence of control actions, but only the actions that fall within the stride interval are retained and implemented. In Figure 2, the retained actions are highlighted. The remaining actions are discarded and shown in light gray.These correspond to future time steps beyond the stride. The action sequences computed by the first and second optimization runs are denoted by $\{x^{1}_{n}\}$ and $\{x^{2}_{n}\}$, respectively.
In the experimental setup considered in this study, the market time resolution is treated as an exogenous parameter and can take values representative of real-world electricity markets, e.g., 15 minutes or 1 hour. The forecast publication interval and the stride of the rolling-horizon optimization are also treated as parameters, and for the purposes of our experiments, they are assigned equal values: each new forecast triggers a new optimization and a corresponding update of the control actions. This choice is practical, as it reflects the fact that no additional information is assumed to become available between forecast updates. A reasonable range is selected for both stride and forecast interval of 1–6 hours, which allows testing a broad spectrum of planning horizons. Very long rolling window strides (e.g., 24 hours) are not suitable for this study. They would set a very large minimum for the size of the optimization window, which is the main parameter that the experimentation aims to explore. In many cases, forecasts are in fact published every 24 hours. But in real deployments, updates triggering re-optimization need not correspond only to new forecasts; they could also arise from real-time measurements of battery state, corrections of forecasts based on updated weather data, or observed production from photovoltaic sources Pinson et al. (2009). This ensures that the methodology remains applicable even when forecasts are published infrequently, as the relevant temporal resolution is defined by the arrival of new actionable information rather than by a fixed publishing schedule.
Finally, while the stride and forecast publication interval are kept fixed within each individual experiment, the length of the optimization window is treated as a free parameter. Identifying the optimization window size that yields the best performance under different data characteristics, forecast uncertainty levels, and battery technologies is the central objective of the methodology presented in this paper.
3.3 Optimization model (MILP formulation)
The battery scheduling problem is formulated as a mixed-integer linear program (MILP) that maximizes total arbitrage revenue over a finite time horizon. The model decides, at each time step $t\in\{1,\dots,T\}$, how much energy to charge/buy and discharge/sell.
Decision variables
At each time step $t$, the following variables are defined:
  • $c_{t}\geq 0$: battery charging power.
  • $d_{t}\geq 0$: battery discharging power.
  • $s_{t}\geq 0$: state of charge (SoC).
  • $g_{t}^{\mathrm{in}}\geq 0$: energy purchased from the grid.
  • $g_{t}^{\mathrm{out}}\geq 0$: energy sold to the grid.
  • $z_{t}\in\{0,1\}$: binary variable enforcing that the battery cannot charge and discharge simultaneously.
  • $y_{t}\in\{0,1\}$: binary variable enforcing that grid import and export cannot occur simultaneously.
Parameters
The system operates with the following parameters:
  • $p_{t}^{\mathrm{buy}}$: electricity purchase price at time $t$.
  • $p_{t}^{\mathrm{sell}}$: electricity selling price at time $t$.
  • $C$: maximum battery capacity.
  • $P_{c}$: maximum charging power.
  • $P_{d}$: maximum discharging power.
  • $\eta$: charging efficiency.
  • $s_{0}$: initial state of charge.
  • $\hat{G}^{\mathrm{in}}$: upper bound on grid import.
  • $\hat{G}^{\mathrm{out}}$: upper bound on grid export.
Objective function
The objective is to maximize trading revenue:
$\max\sum_{t=1}^{T}\left(p_{t}^{\mathrm{sell}}\,g_{t}^{\mathrm{out}}-p_{t}^{\mathrm{buy}}\,g_{t}^{\mathrm{in}}\right).$ (7)
Constraints
(1) Energy balance:
$g_{t}^{\mathrm{in}}+c_{t}=g_{t}^{\mathrm{out}}+d_{t},$ (8)
which enforces that all energy flows are internally consistent.
(2) State of charge dynamics:
$s_{t}=\begin{cases}s_{0}+\eta c_{t}-d_{t},&t=1,\\ s_{t-1}+\eta c_{t}-d_{t},&t>1.\end{cases}$ (9)
(3) Battery operating limits:
$0\leq c_{t}$ $\leq P_{c}z_{t},$ (10)
$0\leq d_{t}$ $\leq P_{d}(1-z_{t}).$ (11)
(4) Grid import/export exclusivity:
$0\leq g_{t}^{\mathrm{in}}$ $\leq\hat{G}^{\mathrm{in}}y_{t},$ (12)
$0\leq g_{t}^{\mathrm{out}}$ $\leq\hat{G}^{\mathrm{out}}(1-y_{t}),$ (13)
(5) State of charge bounds:
$0\leq s_{t}\leq C.$ (14)
Overall, this formulation represents a simple and standard battery arbitrage model in which revenue is obtained by optimally shifting energy across time in response to price fluctuations.
3.3.1 Effective and optimal planning horizons
This section introduces two related but conceptually different notions: the effective planning horizon and the optimal planning horizon under uncertainty. These concepts are illustrated in Figures 3 and 4, respectively.
Effective planning horizon
Figure 3 considers an idealized setting in which the optimization algorithm has access to agnostic (i.e., perfect or uncertainty-free) future values of the underlying signal, such as electricity prices or net demand. On the same timeline, we report the total profit obtained by solving the rolling-horizon optimization problem for increasing sizes of the optimization window.
Figure 3: Illustration of the effective planning horizon under perfect information. The underlying signal (e.g., electricity prices) is shown alongside the profit achieved as a function of optimization window size. Vertical lines mark the myopic horizon (short-sighted planning), the effective planning horizon (window length beyond which additional information does not increase profit), and a redundant horizon (beyond which the window is longer than necessary).
As the planning horizon increases, the achievable profit initially improves, reflecting the algorithm’s ability to exploit additional future information. Beyond a certain horizon length, however, further extending the optimization window does not lead to any additional improvement: the profit curve reaches a plateau. We define the horizon at which this saturation occurs as the effective planning horizon, a concept that is general and applies to any timeseries. Intuitively, it represents the maximum amount of future information that can be effectively exploited in the optimization problem; information beyond this point does not influence optimal decisions.
In mathematical terms, the effective planning horizon can be defined as the horizon length beyond which additional future information does not increase profit. Let $P(H)$ denote the profit obtained by solving the rolling-horizon optimization problem with an optimization window of length $H$ using ground-truth data. Let
$P_{\text{max}}=\max_{H}P(H)$ (15)
be the maximum achievable profit. Then the effective planning horizon $H_{\text{eff}}$ is given by
$H_{\text{eff}}=\min\Big\{H:P(H)\geq(1-\epsilon)\,P_{\text{max}}\Big\},$ (16)
where $\epsilon\ll 1$ is a small tolerance accounting for numerical or rounding effects.
For completeness, a myopic horizon is also marked, corresponding to a short planning window that neglects relevant future effects, and a longer conservative horizon, located beyond the effective planning horizon, where additional look-ahead is redundant. While optimization windows longer than the effective horizon are feasible, they do not yield performance gains and may incur in unnecessary computational costs or slower decision-making processes.
Optimal planning horizon under uncertainty
Figure 4 extends this reasoning to the realistic case in which future values are not known perfectly but are instead provided through forecasts. The figure shows the same underlying ground-truth signal ($GT$), a corresponding forecast ($F$), and the resulting profit as a function of the optimization window size.
Figure 4: Illustration of the optimal planning horizon in the presence of forecast uncertainty. Ground-truth values (GTGT) and forecast values ($F$) are shown along with the resulting profit as a function of optimization window size. Vertical lines indicate the effective planning horizons under perfect information for the ground truth and the forecast, and the optimal horizon, defined as the window length maximizing profit.
In this setting, the profit no longer increases monotonically with the horizon length. Instead, it typically exhibits a maximum: while longer horizons allow the algorithm to anticipate future events, increasing forecast uncertainty eventually degrades decision quality, leading to reduced performance. The optimal planning horizon is the horizon length at which the achieved profit is maximized.
In formal terms, the optimal planning horizon under uncertainty can be defined as the horizon length that maximizes profit when only forecast values are available. Let $\tilde{P}(H)$ denote the profit obtained with an optimization window of length $H$ using forecast data. Then the optimal horizon $H_{\text{opt}}$ is given by
$H_{\text{opt}}=\arg\max_{H}\tilde{P}(H).$ (17)
The difference between the effective and optimal horizons,
$\text{Uncertainty gap}=H_{\text{eff}}-H_{\text{opt}}\geq 0,$ (18)
quantifies the reduction in exploitable information due to forecast uncertainty.
For reference, the effective planning horizon identified under perfect information is also indicated in Figure 4. In general, the effective horizon associated with the ground truth and that associated with the forecast need not coincide. Crucially, when the optimal planning horizon lies to the left of the ground-truth effective planning horizon, it means that the optimization window has been shortened due to forecast uncertainty. In this case, the algorithm deliberately discards potentially useful long-term information because it is too uncertain to be exploited reliably. The interval between the optimal horizon and the effective planning horizon quantifies this loss of exploitable foresight and is referred to here as the uncertainty gap.
Conversely, when the optimal planning horizon coincides with the effective planning horizon, forecast uncertainty does not impose an additional limitation, and no information is discarded relative to the ideal case. By construction, the optimal planning horizon cannot exceed the effective planning horizon defined under perfect information, as information beyond the latter does not contribute to improved performance even in the absence of uncertainty.
3.3.2 Numerical computation of horizon parameters
The experiments systematically evaluate the performance of rolling-horizon optimization across a range of horizon lengths. To identify the effective planning horizon, the optimization window is increased incrementally from zero up to the point at which the achieved revenue saturates, using ground-truth values. To determine the optimal horizon under uncertainty, a similar incremental sweep is performed using forecast values, recording the revenue at each horizon until it reaches a maximum and begins to decrease. This procedure allows quantifying the effective horizon, the optimal horizon, and the uncertainty gap for each dataset and battery configuration.
Revenue under forecast
Each forecasted spot price signal is decomposed into separate consumption (buy) and production (sell) price signals, thereby introducing an explicit spread and avoiding the unrealistic assumption of transacting at a single spot price. Following (Vargunin, 2023, Ch. 3, Eq. (4), p. 19), this transformation is used here in a stylized form, for simplicity, defined as:
[…]
where $p^{\text{spot}}_{t}$ denotes the forecasted spot price at time $t$ in $\mathrm{EUR}$. This formulation ensures that charging and discharging decisions are evaluated under more realistic market conditions by explicitly accounting for transaction asymmetry.
When forecasts are used, revenues are computed by applying the actions determined under the forecast to the actual realized (ground-truth) values. Let $x_{t}(H)$ denote the action taken at time $t$ when using a forecast-based optimization window of length $H$, and let $GT_{t}$ be the realized ground-truth signal (e.g., prices or production). Then the realized revenue is
$\tilde{P}(H)=\sum_{t=1}^{T}x_{t}(H)\,GT_{t}.$ (20)
This ensures that even when the optimization relies on imperfect forecasts, the evaluation measures the revenue that would actually be obtained in practice. Note that for price-based datasets, the ground-truth values are necessary to compute revenue under forecast, whereas for production-based datasets (not studied in this paper), the realized revenue can be computed directly from the forecast actions without reference to additional ground-truth production measurements (although the impact on the charge level of the battery should be considered).
Finally, it is important to emphasize that although one family of datasets used in this study is derived from real mFRR data, the energy arbitrage framework considered here does not reflect the actual operation of the mFRR market. The data are used solely as a reference to capture realistic statistical properties. A simplified energy arbitrage mechanism is adopted to ensure interpretability and enable consistent comparison across experiments.
4 Experiments
This section describes the experimental setup used in this study. It details the synthetic datasets generated for battery scheduling optimization (both ground truth and forecasts) as well as the battery specifications, forecast publication intervals, rolling window strides, and maximum planning horizons tested.
4.1 Datasets
4.1.1 Synthetic actuals
Three families of datasets were generated following the procedure described in the Methods section. One is based on a simple sine wave, to provide a highly interpretable starting point, and the other two are based on real-market datasets (the day-ahead and the mFRR markets). Each dataset spans a two-week period with hourly resolution. This duration was chosen to capture at least one full weekly cycle, which is typically the dominant temporal pattern in energy-related time series like the day-ahead market. While longer periods could potentially provide more comprehensive coverage of seasonal or multi-week effects, computational considerations limit the practical length of the synthetic experiments: the evaluation of multiple forecast horizons across all dataset variants is computationally intensive. Moreover, two weeks of data provides a sufficient visual and conceptual scope for illustrating temporal patterns, forecast performance, and optimal planning horizons, without overwhelming the reader with overly long plots. It is acknowledged, however, that this is a limitation of the current study, and future work could explore longer-term datasets to assess seasonal or inter-week effects more comprehensively.
Sine-wave family
The first family is based on a simple sine wave with a 24-hour daily phase. Three ground truth instances of this family were generated: (i) the undistorted sine wave, (ii) the sine wave distorted by a SARIMA model fitted to a real two-week sample of the day-ahead market, and (iii) the sine wave distorted by a SARIMA model fitted to the mFRR market dataset (a single signal constructed from both mFRR-up and mFRR-down as explained later). Both markets correspond to Estonia. These three instances of the first family (sine-wave family) are illustrated in Figure 5.
The purpose of introducing these distortions is to bridge the gap between a fully controlled synthetic signal and the statistical properties of real-world data. The base sine wave enables the analysis of simple and interpretable behavior under idealized conditions. By injecting SARIMA-based distortions calibrated on real market data, the resulting signals inherit realistic variance and volatility patterns while preserving the underlying daily seasonal structure. This allows isolating the effect of uncertainty and variability without altering the fundamental periodic dynamics. Furthermore, using distortions derived from the day-ahead and mFRR markets ensures consistency with the following dataset families used in this paper, facilitating meaningful comparisons across experiments while maintaining a controlled baseline.
The amplitude of the sine wave was designed such that its estimated maximum potential revenue lies within the range observed for the Fourier-only signal of the day-ahead and mFRR datasets. In this work, the upper bound revenue is approximated directly from the signal as the average total variation per day, defined as
$\hat{R}=\frac{1}{D}\sum_{t=1}^{N-1}\left|p(t+1)-p(t)\right|,$ (21)
where $p(t)$ is the price signal, $N$ the total number of time steps, and $D$ the number of days in the dataset.
Applying Equation 21 to the two-week DA and mFRR datasets yields average values of $37.49~\mathrm{EUR}/day$ and $42.51~\mathrm{EUR}/day$, respectively. Their mean value $40.00~\mathrm{EUR}/day$ was therefore selected as the target estimated maximum potential revenue for the pure sine-wave signal. While it would have been possible to calibrate each distorted sine-wave instance to match the corresponding market-specific value, a single common amplitude was retained across the three instances of the family to preserve consistency and avoid introducing additional degrees of freedom.
Given a target revenue $\hat{R}$, the amplitude $A$ of the sine wave $p(t)=A\sin\left(\frac{2\pi}{T}t\right)$, with period $T=24$, is obtained by normalizing with respect to the total variation of the unit-amplitude sine wave:
$A=\frac{\hat{R}}{\sum_{t=1}^{N-1}\left|\sin\left(\frac{2\pi}{T}(t+1)\right)-\sin\left(\frac{2\pi}{T}t\right)\right|}.$ (22)
Applying Equation 22 with a target revenue of $40.00~\mathrm{EUR}/day$, an amplitude $A=10.00~\mathrm{EUR}$ is obtained.
Figure 5: Synthetic sine-wave dataset illustrating the three ground truth signal variants used in the experiments: the undistorted base sine-wave signal, a day-ahead distorted version, and an mFRR distorted version.
Market-inspired families
The second and third families are inspired by the Estonian day-ahead and mFRR markets, respectively. As mentioned earlier, these datasets are chosen for their relevance, but in this work, they are detached from their market mechanics. Instead, all experiments are executed purely on price arbitrage and without taking into account their corresponding market mechanics. Furthermore, the mFRR dataset used here has been constructed from both sub-markets: mFRR-up and mFRR-down, by combining the two signals according to their activation volumes.
Each family includes three instances formulated as $signal=Fourier+\alpha\cdot SARIMA$, where the Fourier term corresponds to the decomposition of a real two-week sample, using 14 harmonics to capture daily seasonality:
  1. 1.
    A Fourier-only signal: $signal=Fourier+0.0\cdot SARIMA$.
  2. 2.
    A hybrid signal: $signal=Fourier+0.5\cdot SARIMA$
  3. 3.
    A hybrid signal: $signal=Fourier+1.0\cdot SARIMA$
The Fourier-only instance captures the long-term daily pattern, while the hybrid signals incorporate SARIMA components to represent finer-grained variability. The $\alpha$ coefficient controls the SARIMA contribution, allowing evaluation of the impact of variability on battery scheduling. When $\alpha=0$, then the signal is a pure Fourier-only decomposition. It was observed that even a Fourier decomposition with only three harmonics, combined with the SARIMA component, produced visually and statistically satisfactory series for both day-ahead and mFRR markets. The additional harmonics were used to allow testing of the series generated without the SARIMA component, enabling the isolation of long-term patterns from short-term variability. This approach, with a reduced number of Fourier components, provides a foundation for future work: aiming to relate generative data parameters (Fourier + SARIMA parameters) to variables such as optimal horizon length and expected revenue.
It is important to note that when fitting a series with Fourier + SARIMA, the resulting curve closely matches the original data. However, when reproducing a series using only the saved SARIMA parameters, the generated series differs from the original fit: it is qualitatively similar and visually realistic, but the numerical fit is not as close. This occurs because the SARIMA model relies on stochastic residuals during simulation, which are not fully preserved when only the model parameters are stored. The obtained SARIMA parameters for the day-ahead and mFRR ground truth series are presented in Table 1. The mFRR process exhibits lower persistence in both seasonal and short-term autoregressive components, while displaying higher innovation variance, consistent with its more volatile price dynamics compared to the day-ahead market.
Parameter DA mFRR
Non-seasonal order $(p,d,q)$ $(1,0,0)$ $(1,0,0)$
Seasonal order $(P,D,Q,s)$ $(1,0,0,24)$ $(1,0,0,24)$
AR coefficient $ar(L1)$ 0.622 0.467
Seasonal AR coefficient $ar(L24)$ 0.355 0.146
Innovation variance $\sigma^{2}$ 160.55 254.95
Table 1: SARIMA parameters obtained for day-ahead and mFRR price signal generation models.
This study generates synthetic series strictly from the saved SARIMA parameters intentionally, to ensure full reproducibility. Although this reduces the numerical fit (reflected in higher MAE and MSE values), it preserves the essential statistical and visual characteristics of the market series. This approach facilitates the aforementioned future work.
Table 2 summarizes the differences obtained between the synthetic and real datasets.
Figures 6 and 7 show the three instances for the day-ahead and mFRR market-inspired families, respectively.
Dataset Model MAE MSE
Day-ahead Fourier-only 15.01 374.30
Day-ahead Fourier + 0.5 SARIMA 15.83 395.30
Day-ahead Fourier + 1.0 SARIMA 19.31 568.03
mFRR Fourier-only 13.69 319.22
mFRR Fourier + 0.5 SARIMA 15.78 410.93
mFRR Fourier + 1.0 SARIMA 20.70 676.48
Table 2: Accuracy metrics of synthetic series versus real data. Fourier-only series capture the main daily pattern, while the SARIMA component adds variability. MAE: mean absolute error; MSE: mean squared error.
Figure 6: Day-ahead-inspired series illustrating the three reconstruction approaches: the Fourier-only series (capturing daily phase harmonics), the hybrid series with Fourier + 0.5 SARIMA, and the hybrid series with Fourier + 1.0 SARIMA.
Figure 7: mFRR-inspired series illustrating the three reconstruction approaches: the Fourier-only series (capturing daily phase harmonics), the hybrid series with Fourier + 0.5 SARIMA, and the hybrid series with Fourier + 1.0 SARIMA.
Overall, each of the instances (9) from the three families have been chosen or designed within similar price ranges to facilitate comparison of results to a reasonable extent. Prices range roughly between 0 and 150 euros, with eventual negative prices and some peaks close to 200 euros. The sine-wave family was carefully designed to yield profits in a range close to the average between the maximum potential revenues of the day-ahead and mFRR datasets. However, the DA and mFRR signals exhibit slightly different ranges (mFRR generally displaying higher variability). Despite this difference, no additional scaling was applied, seeking to preserve their market characteristics. Both datasets already fall within a comparable range, making further adjustment unnecessary.
4.1.2 Synthetic forecasts
Forecast series are generated following the procedure described in the Methods section. For each synthetic dataset, rolling forecasts with a horizon of 72 hours are produced at intervals of 3 hours; that is, a new 72-hour hourly forecast is issued every 3 hours. All the datasets were also tested for a publishing interval of 6 hours. However, to avoid unnecessary duplication of results, only the 3-hour interval is presented in the results, while the 6-hour case is deferred to the Discussion section. For each of the nine synthetic time-series instances (three families with three variants each), five corresponding forecast series are generated. Every combination is then tested for 6 different batteries, yielding a total of 9 x 5 x 6 = 270 experiments, each of them tested for 45 different planning horizons (8,910 2-week hourly resolution optimization runs).
With regards to the decay of forecast accuracy (increase in the standard deviation of the error distribution), a simple linear approach was chosen to maintain consistency across datasets and ensure interpretability. Based on empirical observations from the real datasets, an exponential increase was deemed unrealistic, as error growth exhibited patterns closer to linear or weakly seasonal behavior. In particular, the day-ahead market showed some intra-day structure, with higher variability during peak hours and lower variability overnight, whereas the mFRR data did not exhibit a clear seasonal pattern. The linear choice provides a simple and transparent representation of forecast degradation while avoiding additional assumptions about intra-day structure.
(a) Day-ahead (DA) dataset: synthetic vs. real MAE.
(b) mFRR dataset: synthetic vs. real MAE.
Figure 8: Mean Absolute Error (MAE) between ground truth and forecast curves for both synthetic and real datasets over a 24h span. Left: day-ahead (DA) series with an uncertainty factor of 1.0; right: mFRR series with an uncertainty factor of 10.0.
As mentioned in the Methods section, the variation between these forecasts is introduced through the width of the error distribution. Specifically, the standard deviation of the forecast error is multiplied by ‘uncertainty factors’ (u.f.) of 0.1, 1, 3, 6, and 10, producing five forecast series with progressively increasing levels of uncertainty. These values have been adjusted (for interpretability) in such a way that the uncertainty factor $u.f.=1.0$ corresponds, roughly, to the uncertainty obtained from real forecasts in the day-ahead market for a span of 24h. While the uncertainty factor $u.f.=10.0$ is comparable to the level uncertainty found in the mFRR market for the same period, as illustrated in Fig. 8.
This design allows evaluating the impact of forecast uncertainty on the variables under study (revenue and optimal planning horizon length). Although varying the autocorrelation structure of the forecast errors could also provide interesting insights, this parameter is kept constant in order to isolate the effect of variance. Figure 9 illustrates a forecast curve example corresponding to the uncertainty factor ($u.f.=1.0$) on the Fourier + 1.0 SARIMA instance of the day-ahead family. Only one forecast curve is shown for clarity, rather than every 3-hour updates.
Figure 9: Example of forecast curve for uncertainty factor (u.f.) = 1.0. The image shows one forecast of the day-ahead series (Fourier + 1.0 SARIMA).
4.1.3 Battery configurations
To assess the impact of storage characteristics on optimal planning horizons, a set of stylized battery configurations is considered, spanning durations from 1 to 24 hours. All configurations are normalized to the same energy capacity (10 MWh equivalent units), while varying power and efficiency to reflect typical trade-offs across technologies (Table 3). The configurations cover short-duration lithium-ion systems (1–4 hours), medium-duration flow batteries (6–8 hours), and long-duration energy storage (LDES) such as hydrogen systems (24 hours), enabling a systematic analysis of how storage duration and efficiency influence optimization performance. All batteries are initialized with a state of charge of 2 MWh.
Table 3: Battery configurations considered in the experiments. All systems have identical energy capacity (10 units), while power and efficiency vary to represent different storage durations and technologies.
Type Duration (h) Power Efficiency Technology
1h Lithium-ion 1 10.00 0.95 Li-ion
2h Lithium-ion 2 5.00 0.95 Li-ion
4h Lithium-ion 4 2.50 0.93 Li-ion
6h Flow 6 1.67 0.80 Flow
8h Flow / LDES 8 1.25 0.78 Flow/LDES
24h LDES 24 0.42 0.40 LDES
5 Results
The results are summarized in Tables 4(a)56 and Figures 101112 with one table and figure per dataset family. The tables report, for each combination of battery configuration and uncertainty factor, the optimal planning horizon. In some cases, relatively long optimal horizons are observed; these typically correspond to plateau-like revenue curves, where the maximum revenue is only marginally higher than that obtained with significantly shorter horizons.
Figures 101112, present the results as 3D needle plots. The x-axis represents the battery cycle time (in hours), the y-axis the uncertainty factor, and the z-axis the optimal planning horizon. The color scale indicates the corresponding revenue. A consistent color scale is used across all datasets to facilitate visual comparison.
Additionally, plots showing the evolution of the revenue as a function of the planning horizon length are also provided in Figures 13141516. In this case, only four datasets have been selected to avoid overwhelming the reader with charts. The datasets chosen are: two sine wave datasets, one with day-ahead SARIMA distorsion and the other one with mFRR distorsion, and the day-ahead + 1.0 SARIMA as well as the mFRR + 1.0 SARIMA datasets, since they are the closest to real-world data. These plots correspond to a publishing interval of rolling forecasts of 3h (as all the experiments presented in this section). However, as indicated earlier, the same plots are provided for a publishing interval of 6h in the Appendix (Figures 24252627) It is also worth noting that the lowest uncertainty factor considered (0.1) yields results that are very close to those obtained under perfect foresight, with differences within approximately 5% of the total revenue.
u.f. 0.1 1 3 6 10
battery
01h 10 4 3 3 3
02h 19 5 3 3 3
04h 18 4 3 3 3
06h 5 3 3 3 3
08h 5 3 3 3 3
24h 22 11 4 4 4
(a) Undistorted sine wave
0.1 1 3 6 10
7 7 6 4 4
9 9 7 6 6
15 11 11 7 6
21 14 10 9 7
22 14 10 10 7
28 12 7 10 5
(b) Sine wave, day-ahead distortion
0.1 1 3 6 10
6 6 5 4 4
13 8 6 5 5
15 25 7 7 8
18 12 9 7 7
18 12 9 7 5
30 12 7 5 5
(c) Sine wave, mFRR distortion
Table 4: Optimal horizon (hours) for different battery types and forecast uncertainty factors (u.f.) of the sine wave dataset family.
u.f. 0.1 1 3 6 10
battery
01h 12 4 3 3 3
02h 12 6 3 3 3
04h 17 6 3 3 3
06h 22 6 3 3 3
08h 28 8 4 3 3
24h 22 9 4 4 4
(a) DA, Fourier-only.
0.1 1 3 6 10
15 18 5 5 4
18 12 5 4 3
24 11 5 3 3
30 5 4 3 3
28 6 4 3 3
30 7 6 4 4
(b) DA, Fourier + 0.5 SARIMA.
0.1 1 3 6 10
9 6 5 4 4
11 9 8 9 5
26 19 8 6 5
29 9 11 5 5
28 10 10 5 5
28 15 9 8 6
(c) DA, Fourier + 1.0 SARIMA.
Table 5: Optimal horizon (hours) for different battery types and forecast uncertainty factors (u.f.) of the day-ahead dataset family.
u.f. 0.1 1 3 6 10
battery
01h 11 6 6 7 7
02h 12 9 5 3 3
04h 15 12 5 3 3
06h 26 7 4 3 3
08h 36 7 4 3 3
24h 29 12 5 3 3
(a) mFRR, Fourier-only.
0.1 1 3 6 10
9 6 6 6 7
12 10 6 8 3
15 10 9 8 3
27 10 7 4 3
29 10 7 4 3
31 12 9 8 3
(b) mFRR, Fourier + 0.5 SARIMA.
0.1 1 3 6 10
10 8 6 7 4
15 11 9 6 8
20 11 13 13 6
30 15 10 9 6
35 15 10 9 6
35 12 9 9 4
(c) mFRR, Fourier + 1.0 SARIMA.
Table 6: Optimal horizon (hours) for different battery types and forecast uncertainty factors (u.f.) of the mFRR dataset family.
(a) Undistorted sine wave.
(b) Sine wave distorted by
DA SARIMA parameters.
(c) Sine wave distorted by
mFRR SARIMA parameters.
Figure 10: Optimal planning horizon and revenue results by battery c-rate and forecast variance factor (uncertainty level) of the sine wave dataset family.
(a) DA, Fourier-only.
(b) DA, Fourier + 0.5 SARIMA.
(c) DA, Fourier + 1.0 SARIMA
Figure 11: Optimal planning horizon and revenue results by battery c-rate and forecast variance factor (uncertainty level) of the day-ahead dataset family.
(a) mFRR, Fourier-only.
(b) mFRR, Fourier + 0.5 SARIMA.
(c) mFRR, Fourier + 1.0 SARIMA.
Figure 12: Optimal planning horizon and revenue results by battery c-rate and forecast variance factor (uncertainty level) of the mFRR dataset family.
Figure 13: Sine wave with day-ahead SARIMA distortion dataset: revenue as a function of planning horizon length for four battery configurations with cycle times of 1h, 2h, 4h, and 6h (as indicated in the plots). Each curve corresponds to a different forecast uncertainty factor. All results are based on rolling forecasts with a 3-hour publishing interval.
Figure 14: Sine wave with mFRR SARIMA distortion dataset: revenue as a function of planning horizon length for four battery configurations with cycle times of 1h, 2h, 4h, and 6h (as indicated in the plots). Each curve corresponds to a different forecast uncertainty factor. All results are based on rolling forecasts with a 3-hour publishing interval.
Figure 15: Day-ahead + 1.0 SARIMA dataset: revenue as a function of planning horizon length for four battery configurations with cycle times of 1h, 2h, 4h, and 6h (as indicated in the plots). Each curve corresponds to a different forecast uncertainty factor. All results are based on rolling forecasts with a 3-hour publishing interval.
Figure 16: mFRR + 1.0 SARIMA dataset: revenue as a function of planning horizon length for four battery configurations with cycle times of 1h, 2h, 4h, and 6h (as indicated in the plots). Each curve corresponds to a different forecast uncertainty factor. All results are based on rolling forecasts with a 3-hour publishing interval.
6 Discussion
On a first inspection of the results, several aspects become apparent and are largely in line with expectations. In particular, datasets with lower variance in the ground truth are consistently associated with substantially lower profitability. This behavior is unsurprising for the cases with $\alpha=0$ (Fourier-only signals), where the reduced variability directly limits arbitrage opportunities. However, this effect is somewhat less intuitive in the $\alpha=0.5$ cases, which still exhibit relatively low revenues despite retaining half of the SARIMA component. For instance, for a 1h battery in the day-ahead market with $\alpha=1$ and an uncertainty factor of 0.1 (i.e., close to oracle conditions), the maximum revenue is approximately 65 EUR. In contrast, under $\alpha=0.5$ and the same uncertainty level, the maximum revenue decreases to approximately 20 EUR, corresponding to a ratio of roughly 20:63 (or 0.31).
For the mFRR family, this ratio is approximately $37:100$ ($0.37$), while the ratio between the two $\alpha$ values was quite higher ($0.50$). See Figures 17, 16 (top-left), and 15 (top-left) for indicative revenue estimates for a 1h battery under $\alpha=0.5$. This discrepancy between ratios can be explained by two reinforcing effects. First, trades are directly scaled by $\alpha$, meaning that revenues are halved when $\alpha=0.5$. Second, some trades are effectively lost because their price differentials fall within the bid–ask spread, making them non-executable. This mechanism also helps explain why the ratio is lower in the day-ahead family, which exhibits a less spiky and more contained variance range (as captured by the SARIMA model). In this case, a larger proportion of signal deltas falls within the spread, increasing the number of missed trading opportunities. These observations have an important practical implication: when forecasting price signals, particularly when their variability is close to the bid–ask spread, predictive performance should prioritise correctly classifying whether price deltas exceed the spread threshold, rather than focusing solely on accurately reproducing continuous price levels.
(a) Day-ahead, Fourier + 0.5 SARIMA.
(b) mFRR, Fourier + 0.5 SARIMA.
Figure 17: Day-ahead (left) and mFRR (right) Fourier + 0.5 SARIMA datasets: revenue as a function of planning horizon length for a battery configuration with cycle time of 1h. Each curve corresponds to a different forecast uncertainty factor. All results are based on rolling forecasts with a 3-hour publishing interval.
When these same values are compared for larger uncertainty factors, the ratios decrease even further. For example, for an uncertainty factor of 1.0 (corresponding to a DA-like forecast noise level under the benchmark calibration), the ratios become approximately $12:58$ (or $0.30$) for the day-ahead family, and $30:97$ (or $0.20$) for the mFRR family (see Figures 17, 16 top-left, and 15 top-left). This clearly indicates that uncertainty has a more detrimental effect on price signals with narrower variance bands. This behavior can be explained, at least in part, by the construction of the synthetic forecasting framework used in this study. Since forecast errors are drawn from a fixed error distribution with an additional autocorrelation component, signals with higher intrinsic variance produce larger price deltas that are less likely to be reversed by noise. In contrast, for narrower variance signals, small perturbations are more likely to flip the sign of price deltas, leading to incorrect trading decisions and thus higher losses.
This observation has important implications for predictive modelling. In particular, it suggests that forecasting performance should be evaluated not only in terms of point accuracy, but also in terms of its ability to preserve the correct sign and magnitude of price differentials relative to the trading spread. Finally, it is worth noting that in these cases with narrow variance bands, revenue losses due to uncertainty can be substantial even for very fast batteries (1h cycle), as illustrated in Figure 18.
(a) Day-ahead, Fourier + 0.5 SARIMA.
(b) mFRR, Fourier + 0.5 SARIMA.
Figure 18: Day-ahead (left) and mFRR (right) Fourier + 0.5 SARIMA datasets: revenue as a function of planning horizon length for a battery configuration with a cycle time of 1h and an uncertainty factor of 3.0. The revenue drop due to uncertainty is defined as the relative decrease from the maximum revenue.
As expected, when running the optimizations using Fourier-only ground truth signals (i.e., without SARIMA components), the resulting revenues are significantly lower. For the lowest uncertainty factor ($u.f.=0.1$, close to oracle conditions), the revenues across the three dataset families are only approximately related to the values predicted by Equation 21, but with noticeable deviations, as shown in Table 7. In particular, while the analytical expression provides a reasonable upper-bound scaling based on signal variation, the observed revenues are not strictly proportional to these estimates, and the relative ranking between datasets is not preserved. According to the estimates, the day-ahead dataset would be expected to yield the lowest revenue under Fourier-only ground truth at the lowest uncertainty level. However, in the simulations, the regular sine wave dataset instead produces slightly lower returns. This indicates that, while the overall dependence on signal variation is partially captured by the proxy model, additional structural properties of the signal—such as the distribution of steep transitions and local extrema—also play a role in determining profitability.
When SARIMA-based distortions are introduced on top of the regular sine wave function (i.e., day-ahead and mFRR SARIMA variants), revenues can exceed those of the corresponding full SARIMA datasets (with $\alpha=1.0$). In particular, the maximum revenues for the day-ahead + 1.0 SARIMA case oscillate around approximately 64 EUR/MWh/day, whereas the sine wave dataset with day-ahead distortion reaches values close to 100 EUR/MWh/day (ratio $64:97$). For the mFRR case, this ratio is considerably more balanced (approximately $101:107$). While further experimentation is required to fully explain these effects, a plausible cause lies in the interaction between the Fourier signal phase structure and the localized spikes and dips introduced by the SARIMA components, which may lead to constructive or destructive alignment effects in trading opportunities.
Dataset Upper bound estimation (EUR/MWh/day) Obtained revenue (EUR/MWh/day, $u.f.=0.1$)
Sine wave 40.0 $\sim 6.1$
Day-ahead 37.5 $\sim 7.0$
mFRR 42.5 $\sim 10.1$
Table 7: Comparison between estimated upper-bound revenues and obtained revenues under low uncertainty ($u.f.=0.1$) for different datasets. The upper bound is derived from the total variation of the price signal, while obtained revenues correspond to optimized dispatch results.
6.1 Optimal planning horizon
As a general trend, it is expected that slower batteries tend to require longer planning horizons. In the results of this study, this is mostly true when there is low uncertainty ($u.f.\leq 3.0$), with some exceptions (e.g.: Figure 10(a)). But with higher uncertainties, slow batteries cannot exploit longer horizons without degrading their revenues, as is clearly visible in Figures 101112. Slow batteries with cycles of 8h and 24h can present very long optimal horizon under the lowest uncertainty factor tested (0.1). Specifically, they can reach beyond 30h, especially in the mFRR dataset family. This behavior is consistent across the different 3D plots presented in this study and reflects the limited operational flexibility of slower storage systems. Because these batteries take longer to complete charge–discharge cycles, their scheduling decisions are naturally influenced by uncertain conditions further into the future.
Fast batteries, in contrast, can maximize their revenues with much shorter horizons, typically between 8h and 12h for 1h and 2h cycle batteries. However, when uncertainty is in the range of 1.0, these optimal horizons shrink significantly to a range closer to 4h to 9h (Tables 4(a)56). They usually do so without any revenue drop after the optimal horizon. This last aspect is important because an accurate selection of the planning horizon within the level of uncertainty of the day-ahead market may require much less computation time and cost than otherwise anticipated, without loss in revenue. However, with fast batteries, increasing uncertainty seems to give rise to a regime characterised by a sharp and locally unstable optimum. In these cases, neighbouring horizon values yield considerably less revenue, without a smooth degradation transition, as shown in Figures 18(b) and 19(b). The challenge in these cases is that attempting to find the optimal horizon may lead to very short horizon values. These often fall into the ramping-up section of the curve, which is highly sensitive and can result in substantially lower returns than the optimal.
Another aspect worth noting is that datasets with higher variability in the ground truth signal present longer optimal horizons. For example, Fourier-only signals in all three families have the shortest horizons, followed by Fourier + 0.5 SARIMA signals, and then the full SARIMA signals. These last ones present the longest optimal horizons on average. However, as discussed at the beginning of this Discussion section, there may be an effect at play regarding the actual or effective uncertainty present in these datasets. When the signal variability is low, the synthetic scheme implemented for the generation of forecast values is more likely to flip price deltas than in signals with higher variability, causing a much larger impact on revenue loss. With this information at hand, it may well be that the real reason why the optimal horizons are shorter in datasets with narrower variance bands is that they experience higher effective uncertainty. This issue is a limitation of the present study that should be addressed in future work.
It is also important to note that the optimal planning horizon can exhibit strong sensitivity, even under apparently similar conditions. In particular, the revenue curves often contain extended plateau regions around the maximum, such that a substantially shorter horizon already achieves revenues that are not significantly lower than the optimum (often within less than 0.01 EUR), as shown in Figures 19(a) and 20(b). As a result, the exact location of the optimal horizon can vary considerably without implying meaningful differences in performance. This observation could motivate the definition of a ‘business’ optimal and effective horizon in future studies.
The choice of an optimal planning horizon is important for maximizing revenue. In practice, if the optimal horizon can be predicted or estimated a priori based on properties or features of the incoming data (before solving the optimization problem), it would be possible to adaptively select the horizon in a dynamic manner for each optimization run. Such a dynamic adjustment would likely improve revenues further, since in this study the horizon is kept static for interpretability. In future work, more flexible horizon strategies should be explored. These include soft horizon formulations, abrupt horizon termination combined with learned terminal state-of-charge constraints (noting that in this study no terminal constraint is imposed, or equivalently a zero final state-of-charge is enforced), as well as uncertainty-dependent decay functions that gradually reduce the influence of future time steps. Such approaches would correspond to a form of soft robust optimization over the planning horizon.
Finally, beyond revenue maximization, the selection of an appropriate planning horizon can also reduce computational cost and enable faster decision-making and trading. This becomes particularly relevant when the optimization problem is computationally expensive and market time windows are short. It can also become crucial in settings where batteries are physically distributed at the retail level while optimization is performed centrally by a provider. In these cases, frequent rolling-horizon optimization may lead to non-negligible computational overhead, potentially affecting the economic viability of retail-scale products.
(a) mFRR, Fourier + 1.0 SARIMA.
(b) Day-ahead, Fourier + 1.0 SARIMA.
Figure 19: Illustration of plateau effects in revenue as a function of planning horizon: (a) mFRR market, and (b) Day-ahead market, both under Fourier + 1 SARIMA models. In (a), the revenue curve exhibits a broad plateau around the maximum, indicating that shorter horizons can achieve near-optimal performance. This suggests that a ‘business’ optimal horizon (trading off marginal revenue gains against operational simplicity) would be approximately 9h or 10h instead of 15h. In (b), this plateau is formed after the optimal horizon affecting the effective horizon, which can be visually inferred to be around 8h instead of 34h.
6.2 Effective planning horizon
The experiments consistently show the presence of an effective planning horizon, as formulated theoretically in the Methods section. This horizon represents the point beyond which additional forecast information does not lead to further improvements in revenue. In practical terms, it corresponds to the horizon length after which the battery scheduling problem becomes largely insensitive to additional future information, as the revenue curve tends to plateau. Therefore, it provides a measure of the amount of future information that effectively influences the scheduling algorithm, and it can be useful in itself for understanding the impact of different data profiles and uncertainty levels on planning. In this sense, the effective horizon appears to be primarily driven by structural properties of the underlying time series, given a fixed battery configuration, rather than hinging heavily on the level of uncertainty. This aspect makes it a potentially robust descriptive metric. Along this line, it may also be interesting in future work to formally prove the existence of this horizon, and to characterise the mathematical conditions under which it emerges. Additionally, since it is mostly a property of the interaction between a given time series and the battery system, rather than purely of the interplay between uncertainty and ground truth structure, it could also be valuable to explore analytical expressions for its length.
Although this paper has focused more systematically on the optimal horizon due to its greater practical implications, some anecdotal observations on the effective horizon are also reported here. A dedicated study focusing specifically on the effective horizon would be valuable for future work.
The effective planning horizon observed in the experiments is, in many cases, longer than might be intuitively expected, even for relatively fast batteries. For example, in the case of 4-hour cycle batteries (Figure 20), the effective horizon can reach values beyond 24 hours under low to moderate uncertainty ($u.f.=0.1$ and $u.f.=3.0$). Even for faster batteries, such as those with a 2-hour cycle, the effective horizon may still extend up to 10 to 12 hours in some cases (Figures 131416). This suggests that, in practice, operational decisions may still depend on information relatively far into the future, even when the physical flexibility of the battery is high. In general, fast batteries (1h and 2h cycles) tend to exhibit a relatively clear transition to a flat revenue region shortly after the peak defined by the optimal horizon. These cases provide clear examples of the effective horizon. Batteries with a 1h cycle typically exhibit effective horizons within 5 to 8 hours, while 2h batteries fall in the range of 7 to 12 hours. However, for batteries with 4h cycles and beyond, this flat revenue region becomes less pronounced. In these cases, typical effective horizons range from 16 to 24 hours for 4h batteries, while 6h batteries range from 20 to 36 hours and beyond.
One aspect that becomes apparent is that uncertainty does not exhibit a clear correlation with longer effective horizons, as hinted earlier. There is one exception to this observation, namely when very large uncertainty is present ($u.f.=10$). In this case, small revenue fluctuations are observed for very fast batteries after extended stable revenue regions, pushing the effective horizon beyond 30 hours, as shown in Figure 21. This is likely due to the way synthetic forecasts are generated in this study, which can produce very wide deviations when uncertainty is high. Regardless, it is theoretically interesting that batteries with such short cycle times may still be affected by information so far into the future under such conditions. An additional implication of these observations is that, given the presence of small revenue fluctuations after extended constant-revenue regions exceeding 20 hours (Figures 19(b)21(a)), it is possible that similar effects exist in other experiments but remain unobserved (and unreported) due to the 36-hour horizon limit of the current experimental setup.
Another aspect worth noting is that the horizon appears to increase in a nonlinear manner with the battery’s cycle time. This is not only reflected in the magnitude of the horizon, but also in its variability range. In particular, long-cycle batteries exhibit effective horizons ranging between approximately 20 and 36 hours, while fast-cycle batteries are limited to roughly 5 to 8 hours. In addition, no clear correlation is observed between dataset families, suggesting that this horizon may be more sensitive to specific signal realizations rather than to structural characteristics alone. This could potentially complicate the derivation of analytical expressions for the effective horizon length.
(a) Sine wave + mFRR SARIMA distortion.
(b) mFRR, Fourier + 1.0 SARIMA.
Figure 20: Two examples of effective horizons in 4h batteries, (a) Sine wave with mFRR distortion and u.f. = 3.0, and (b) mFRR with Fourier + 1.0 SARIMA and u.f. = 0.1. In both cases, despite the wide difference in uncertainty factors (3.0 vs. 0.1), quite long effective planning horizons are observed in 4h batteries, reaching 26h in (a) and 29h in (b).
(a) Sine wave + day-ahead SARIMA distortion.
(b) Day-ahead, Fourier + 1.0 SARIMA.
Figure 21: Two examples of unusually long effective planning horizons for fast batteries under high uncertainty conditions ($u.f.=10$): (a) 1h battery and (b) 2h battery. In (a), the effective horizon exceeds the study bound (36 h), while in (b) it remains within but still significantly extended. These effects appear as isolated irregularities in the revenue curves and are not representative in frequency, yet they highlight that, under high uncertainty, even fast-response assets may exhibit unexpectedly long effective planning horizons.
In more practical terms, the effective horizon could serve as an upper bound for the planning horizon that should be considered in the optimization of battery schedules, thereby reducing computational time and costs. This is particularly relevant when no significant drop in revenue is observed after the optimal horizon, or in other words, when the optimal and effective horizons coincide. When a drop in revenue is present, the optimal horizon should be targeted for obvious reasons; however, even in this case, the effective horizon can still provide a useful upper bound for the search space. Lastly, as already discussed for the optimal horizon, it should be noted that the identification of the effective horizon is sensitive to small horizontal fluctuations in the revenue curves. Once the revenue curve approaches its plateau, small variations in value may artificially extend the apparent location of the effective horizon. As a result, the measured horizon may appear longer than what would be operationally meaningful, since several neighbouring horizon values may yield nearly identical outcomes (see Figure 19).
6.3 Revenue degradation and uncertainty gap
The results also reveal the presence of what was described earlier as an uncertainty gap between the optimal planning horizon and the effective planning horizon. This region corresponds to the interval where additional forecast information exists but does not translate into improved operational outcomes due to forecast uncertainty. These losses are measured as the difference between the maximum profit (achieved at the optimal horizon) and the minimum profit observed beyond this point, expressed as a percentage, as defined in Equation (23):
$\mathrm{Loss}~[\%]=\frac{\max_{h}\,R(h)-\min_{h\geq h^{*}}\,R(h)}{\max_{h}\,R(h)}\times 100$ (23)
The uncertainty gap appears in most scenarios, and it tends to become more pronounced as forecast uncertainty increases, as expected. Although the relative losses observed in the experiments can be inferred visually from Figures 131415, and 16 for experiments with rolling forecasts published every 3 hours, and from Figures 242526, and 27 in the Appendix for 6-hour publication intervals, some representative numerical examples of revenue loss are also provided below for reference.
In many cases, the associated reduction in revenue is relatively modest, suggesting that operating slightly beyond the optimal horizon may still lead to near-optimal performance. In particular, for $u.f.\leq 1.0$, revenue losses are generally not substantial, except for batteries with cycle times $\geq 6h$. Another exception is observed in datasets with lower ground-truth signal variability. For example, in the day-ahead and mFRR datasets with $\alpha=0.5$, substantial losses are reported for 4h batteries (30.0%) and smaller losses (2.6%) for 2h batteries even under low uncertainty ($u.f.=1.0$), as shown in Figure 22.
(a) mFRR dataset, Fourier + 0.5 SARIMA.
(b) Day-ahead dataset, Fourier + 0.5 SARIMA.
Figure 22: Revenue as a function of planning horizon under low uncertainty ($u.f.=1.0$): (a) mFRR dataset for a 2h battery, and (b) day-ahead dataset for a 4h battery, both using Fourier + 0.5 SARIMA. In both cases, noticeable declines in revenue are observed beyond the optimal horizon, leading to moderate losses for the 2h battery and more sizeable losses for the 4h battery.
However, for $u.f.\geq 3.0$, revenue degradation after the optimal horizon starts to become substantial. For illustration purposes, two examples of sizeable losses for fast batteries (cycle times of 1h and 2h) are presented in Figure 23. In these cases, losses of 4.2% and 6.3% are observed, which can represent significant absolute values in large-scale battery deployments. For a representative 50 MWh system, such percentages may correspond to losses on the order of several thousand euros per month. These observations are also relevant because they highlight that forecast uncertainty can affect operational performance even for highly flexible storage systems. In other words, fast-response batteries are not fully immune to the adverse effects of uncertainty in the forecasting horizon.
(a) Sine wave + mFRR SARIMA distortion.
(b) Sine wave + mFRR SARIMA distortion.
Figure 23: Revenue as a function of planning horizon under moderate uncertainty ($u.f.=3.0$): (a) sine wave with mFRR distortion dataset for a 2h battery, and (b) the same dataset for a 4h battery. In both cases, noticeable declines in revenue are observed beyond the optimal horizon, leading to moderate losses for both batteries.
In practice, the performance of very fast batteries may also be constrained by operational limitations that prevent them from fully exploiting their theoretical flexibility. For example, operators may impose limits on the number of daily charge–discharge cycles in order to reduce battery degradation and extend system lifetime. Similarly, technical or regulatory constraints may restrict the maximum volume of energy that can be injected into the grid within a given time interval. These types of constraints may partially offset the theoretical advantages associated with very high c-rates. Consequently, theoretically ‘fast’ battery systems may become critically exposed to uncertainty under such conditions, which makes the observations discussed on revenue losses even more relevant.
Beyond operational constraints, market design and capacity remuneration mechanisms can further influence storage technology selection. In several electricity markets, longer-duration storage is implicitly incentivized by tying capacity credit and market participation to a minimum discharge duration. This effectively strengthens the relative competitiveness of multi-hour battery systems in investment decisions. As a result, developer preferences in some regions increasingly align with 4h lithium-ion configurations, which are frequently selected in capacity-market-backed projects. As noted recently (2023) by Denholm et al. Denholm et al. (2023), “Several regions, including CAISO, MISO, NYISO, and SPP have established 4 hours as the minimum duration required to receive full capacity credit” (p. 6). This is important because of their higher exposure to forecast uncertainty and revenue degradation, as reflected in the results of this study.
Building on the interaction between operational constraints and market-driven incentives discussed above, there are several additional factors that may further accentuate the revenue losses associated with sub-optimal planning horizons.
  1. 1.
    First, it should be noted that in this study the planning horizon is fixed over the entire simulation period (2 weeks). The longer-term objective of the analysis is, however, to characterise optimal horizon lengths as a function of data properties and battery design, with the aim of enabling dynamic rather than static horizon selection. If such adaptive horizon selection were implemented, it is expected that optimal profits would be higher than those reported here. Consequently, the losses obtained under a fixed-horizon assumption can be interpreted as a conservative baseline (actual losses would be even worse). This comparison is particularly relevant given that a significant share of current industry implementations still rely on fixed-horizon optimization schemes.
  2. 2.
    Second, the reported losses within the uncertainty gap correspond to reductions in top-line revenue. In realistic economic settings, where operating expenditure (OPEX) and capital expenditure (CAPEX) must be recovered from these revenues, even moderate percentage losses can have a large impact on net profitability. This especially true when profit margins are tight. For illustration, one may consider a simple case where a battery generates 100 EUR of daily revenue under optimal conditions. A 5% reduction corresponds to 95 EUR of revenue. Assuming fixed daily costs of 80 EUR, profit would decrease from 20 EUR to 15 EUR, representing a 25% reduction in profit despite only a 5% drop in revenue.
  3. 3.
    Finally, when multiple correlated signals are used in the optimization process (e.g. production, consumption, and price forecasts), uncertainties may compound. While each individual signal may be associated with an uncertainty level $x$, their joint use in a coupled optimization problem can lead to an effective amplification of uncertainty due to error propagation across interacting inputs. In general, the resulting uncertainty depends on the correlation structure between signals, and may scale between $O(x)$ in the case of independent errors and up to $O(nx)$ in the worst case of fully aligned errors across $n$ signals (which is common in energy markets). This effect can further increase the sensitivity of the scheduling solution to forecast errors, particularly in multi-signal or multi-market optimization settings.
For all these reasons, the results of this study highlight the importance of carefully selecting the planning horizon, as sub-optimal choices can lead to non-negligible revenue losses, particularly under realistic levels of uncertainty and operational constraints.
7 Conclusions and future work
This work has introduced a methodological framework for analysing planning horizon selection in rolling-horizon optimization under forecast uncertainty, with the aim of providing interpretable and practically relevant insights for energy system operation. A central design choice of the framework is the explicit parameterization of the information structure, as well as other aspects like forecast publication intervals or the rolling-horizon strides. By treating these elements as exogenous and configurable parameters rather than fixed assumptions, the approach remains applicable across a wide range of operational settings and lays the groundwork for future models capable of inferring optimal horizons directly from time-series parameters.
Within this framework, the notions of optimal and effective planning horizons are utilized as complementary concepts for understanding the value of future information. The effective planning horizon captures the maximum extent to which future information can influence scheduling decisions, while the optimal planning horizon reflects the horizon length that maximises performance under forecast uncertainty. The gap between these two horizons provides a natural measure of how uncertainty limits the exploitable value of information, and offers a principled basis for reasoning about horizon selection and truncation in rolling-horizon and model predictive control schemes.
The study has examined how battery characteristics, signal structure, and forecast uncertainty jointly influence the optimal planning horizon and resulting revenue in energy storage optimization problems. Across a wide range of synthetic market-like datasets and battery configurations, several consistent patterns emerge that provide insight into both the structure of optimal decision-making and its sensitivity to uncertainty.
  • Behavior of the optimal planning horizon: The results show that the optimal planning horizon is not only shifted by battery characteristics, but also exhibits strong dependence on the statistical structure of the input signal and, most importantly, on forecast uncertainty. Fast batteries with 1–2 hour cycles consistently achieve maximum performance at relatively short planning horizons, typically in the range of 4–8 hours depending on the uncertainty (with stronger uncertainty, shorter horizons). In contrast, slower batteries with 4–6 hour cycles and above require much longer horizons. Under very low uncertainty, they can move between 12 and 24 hours. While under strong uncertainty, this bracket shrinks significantly (5 to 12 hours).
    Long cycle batteries also exhibit substantially greater variability in the location of the optimum, compared to those with short cycles. As a consequence, optimal horizons may result harder to estimate or generalise for long-duration storage systems. In contrast, fast batteries display much more stable and concentrated optimal horizon regions, suggesting a more predictable operational structure.
  • Effect of uncertainty on optimal horizons: Across all battery types, increasing forecast uncertainty leads to a systematic reduction in optimal planning horizons. For a given battery c-rate, the optimal horizon decreases more or less linearly with the increase in uncertainty factor. This reflects the diminishing value of long-term forecast information when predictions become less reliable. In other words, under higher uncertainty, the optimization cannot effectively exploit distant future information, and therefore shorter horizons become optimal.
    Despite this general contraction, an important relative pattern persists: slower batteries continue to require longer horizons than fast batteries even under high uncertainty. This indicates that physical flexibility constraints remain a dominant factor in determining horizon length, even when forecast quality deteriorates significantly.
  • Implications for horizon selection in battery scheduling: The strong sensitivity of the optimal horizon has two important implications for practical optimization. First, it confirms that the planning horizon is a critical decision variable with a strong impact in revenue. Second, because the optimal horizon is highly sensitive to both uncertainty and signal structure, different optimization runs will not necessarily correspond to identical optimal horizons.
    These observations naturally motivate adaptive strategies in which the planning horizon is dynamically adjusted for each optimization run based on current signal and forecast conditions. Such an approach would likely improve performance, particularly in regimes where the revenue surface is locally unstable or exhibits sharp optima. In this sense, horizon selection itself becomes a control problem, rather than a static parameter choice.
  • Existence of an effective planning horizon. The empirical results consistently indicate the presence of an effective planning horizon, defined as the point beyond which additional forecast information does not lead to further changes in revenue. In practical terms, this corresponds to a plateau in the revenue curve, where extending the horizon yields negligible gains. This observation suggests that the effective horizon is a structural property of the interaction between the signal and the battery system, and it motivates future work aimed at formally establishing its existence and characterising it mathematically.
  • Revenue degradation and the uncertainty gap: Revenue degradation beyond the optimal horizon becomes increasingly pronounced as uncertainty increases. For batteries with cycle times up to 4 hours, this degradation becomes particularly acute at uncertainty factors around $u.f.=3$, which represents a regime of moderately high uncertainty beyond typical day-ahead conditions ($u.f.=1$) and below highly volatile markets such as mFRR ($u.f.=10$).
    Batteries with longer cycle times (6 hours and above) already show sensitivity to lower uncertainty regimes, including those representative of day-ahead markets. This suggests that long-duration storage systems are inherently more exposed to forecast errors, even under relatively benign market conditions.
    It is also important to note that uncertainty may be amplified in multi-signal optimization settings. When multiple correlated forecasts (e.g., price, demand, and generation) are jointly used, individual uncertainty levels may compound through error propagation, potentially leading to significantly higher effective uncertainty than suggested by any single input.
  • Sensitivity to small price differentials and structural effects: The results further indicate that when price differentials are small relative to the bid–ask spread, forecasting errors can have an amplified impact on trading decisions. In such regimes, even minor prediction errors can flip the sign of price deltas or incorrectly classify trades as executable or non-executable.
    This effect is particularly relevant in low-volatility regimes, such as certain periods of the day-ahead market, where price signals are smoother and less dispersed. In these cases, the study shows that even medium-cycle batteries (e.g., 4-hour systems) can experience substantial revenue degradation, in some cases reaching losses on the order of 30%.
    This finding is especially important given the increasing policy emphasis on 4-hour storage systems in many regulatory frameworks. It suggests that such systems may be more exposed to forecast structure than commonly assumed, particularly in low-volatility environments.
  • Implications for forecast evaluation and trading-aware metrics: Finally, the results suggest that traditional forecast evaluation metrics based on point-wise accuracy are insufficient for capturing the operational value of predictive models in energy trading contexts. Instead, forecasting performance should be assessed in terms of its ability to correctly identify trading-relevant events.
    In particular, greater emphasis should be placed on correctly classifying whether price differentials exceed the trading threshold defined by the bid–ask spread, as well as whether the direction (sign) of price changes is correctly predicted. This framing shifts the focus from continuous signal reconstruction to decision-relevant classification of profitable trading opportunities.
Overall, these findings highlight that optimal scheduling in battery systems is governed by a complex interaction between physical constraints, signal structure, and forecast uncertainty. Accounting for these interactions is essential for designing robust and economically efficient storage operation strategies.
7.1 Future work
While the present study provides a controlled and interpretable analysis of the interaction between signal structure, forecast uncertainty, and planning horizon selection, two important limitations should be prioritised in future work. First, the analysis relies on parametrically generated datasets, which, although useful for isolating specific effects, do not fully capture the complexity of real-world market signals. A systematic comparison with real datasets is therefore necessary to assess how well the proposed parametrization and resulting insights generalise to practical settings. Second, the experiments are conducted over relatively short time horizons and with a single instantiation of the stochastic forecasting process. Extending the analysis to longer time periods would allow the inclusion of seasonal effects (e.g., summer and winter dynamics), structural changes, and rare events. Also, evaluating multiple stochastic realizations (i.e., different random seeds) would enable a more robust statistical characterization of the results. Together, these extensions would strengthen the empirical validity of the findings and provide a more comprehensive assessment of horizon selection under realistic operating conditions.
In addition, several other directions emerge from this study that call for further research:
  • Prediction of optimal horizon from data features: A natural continuation of this work, as mentioned several times throughout the paper, would be to investigate whether the optimal (or near-optimal) horizon can be predicted directly from observable properties of the input signals (e.g., variance, frequency content, or uncertainty metrics) prior to solving the optimization problem. This would improve revenues, enable faster decision-making and reduce computational overhead.
  • Soft horizon formulations: The current framework relies on a hard truncation of the planning horizon. Future research could explore ‘soft’ horizon approaches, including learned terminal state constraints (e.g., terminal state-of-charge), or decay functions that progressively reduce the influence of distant forecasts. Such methods could provide a more robust treatment of uncertainty in long horizons.
  • Theoretical characterization of the effective horizon: While the existence of an effective planning horizon is empirically observed, its theoretical properties remain unexplored. Future work could aim to formally establish conditions under which an effective horizon exists, and to derive analytical expressions or bounds for its length as a function of signal and system parameters.
  • Improved modelling of forecast uncertainty: The synthetic forecasts used in this study allow for large deviations under high uncertainty levels, which may influence some observed effects (e.g., extended effective horizons at extreme uncertainty). Future work should consider more realistic uncertainty models, including calibrated probabilistic forecasts and empirically derived error distributions.
  • Extension to uncertainty-aware optimization frameworks: While this study intentionally relies on deterministic optimization in order to isolate and interpret the effects of forecast uncertainty, an important extension would be to repeat the analysis under uncertainty-aware formulations, such as stochastic or robust optimization. In particular, it would be valuable to investigate how the optimal and effective planning horizons behave when uncertainty is explicitly modelled within the optimization problem.
  • Multi-signal and multi-market interactions: This study highlights the potential for uncertainty amplification when multiple correlated signals are used jointly. Future research should systematically investigate this effect in coupled optimization settings (e.g., co-optimization of energy and ancillary services), including the role of correlation structures and error propagation mechanisms.
  • Impact of operational constraints: The analysis assumes idealised battery operation. Incorporating more detailed operational constraints, could provide a more realistic assessment of the interaction between planning horizon and profitability. Such aspects may include cycle limits, degradation costs, and efficiency variations.
  • Extension to broader market mechanics: Given that market rules (e.g., bid-ask spreads, capacity requirements, and settlement intervals) influence the results, it would be valuable to extend the analysis to a wider range of real market structures. This includes evaluating how different regulatory frameworks affect optimal horizon selection and sensitivity to uncertainty.
Taken together, these directions highlight that the present study should be viewed as a controlled and intentionally simplified framework aimed at isolating key structural effects in a transparent manner. While this design choice enables interpretability and clear attribution of observed effects, it also necessarily takes away several dimensions of real-world complexity. Addressing the outlined limitations will therefore be essential for validating the robustness of the proposed insights and for translating them into operational decision-making tools in realistic energy market environments.
Author contributions
J. de-Miguel-Rodriguez led the study, including the initial ideation, methodological development, experimental design, implementation, and manuscript preparation. A. Vargunin contributed to the ideation and conceptualization of the approach and implemented supporting scripts used in the experimental evaluation. B.R. Raudne developed the synthetic data generation framework and produced several visualizations used in the analysis. D. Solis-Martin implemented the experimentation engine used to run and evaluate the proposed methods and provided the ‘real’ forecasts used in the study. Y. Mykhailenko contributed to data analysis and interpretation of the experimental results. K. Oja contributed to project supervision, funding acquisition, and overall support of the research activities. All authors have reviewed and approved the final manuscript.